### v3

# GL1-22-11-24-2-supplement

Abstract
In this paper, we develop a Difference-in-Differences model for discrete, ordered outcomes, building upon elements from a continuous Changes-in-Changes model. We focus on outcomes derived from self-reported survey data eliciting socially undesirable, illegal, or stigmatized behaviors like tax evasion or substance abuse, where too many “false zeros”, or more broadly, underreporting are likely. We start by providing a characterization for parallel trends within a general threshold-crossing model. We then propose a partial and point identification framework for different distributional treatment effects when the outcome is subject to underreporting. Applying our methodology, we investigate the impact of recreational marijuana legalization for adults in several U.S. states on the short-term consumption behavior of 8th-grade high-school students. The results indicate small, but significant increases in consumption probabilities at each level. These effects are further amplified upon accounting for misreporting.
JEL Codes: C21, C25, C31, C51
Keywords: Partially Observable Outcomes, Changes-in-Changes, Identification, Measurement Error
1 Introduction
The Difference-in-differences (DiD) method is widely recognized in the econometric literature as a quasi-experimental technique for identifying causal mean parameters. While this method has been central in understanding the average effects of reforms or policy changes, there has been a growing interest in capturing the full spectrum of counterfactual outcomes as policy interventions may have multifaceted impacts across the entire distribution. Consequently, quasi-experimental methods exploiting pre- and post-treatment data have emerged over the past two decades to identify and estimate such distributional treatment effects (see e.g., Roth and Sant’Anna, 2023, for an overview). A prominent example of this more comprehensive approach is the Changes-in-Changes (CiC) model developed by Athey and Imbens (2006).
This paper presents a CiC model for discrete, ordered outcomes that are potentially underreported, nesting the “misreporting-to-zero” or “false-zero” scenario as a special case. Underreporting is prevalent in survey data on discrete ordered outcomes of count, ordinal, or binary nature, especially when socially undesirable, illegal, or stigmatized behaviors and actions are elicited, posing significant challenges for both model identification as well as estimation and inference. Common examples include the individual consumption of “social bads” such as marijuana (cannabis) and alcohol, and sensitive issues like abortion or domestic violence incidents. [^1] For example, Brown et al. (2018) found that the actual incidence of marijuana consumption in an Australian survey was nearly double the reported figure (23% vs. 12.2%), while Caudill and Mixon (2005) identified about 20 percentage points underreporting in cheating behavior among undergraduate students in a U.S. survey. See also Nguimkeu et al. (2019) for other examples.
The paper makes several contributions to the literature. Firstly, we consider a general ordered choice model with underreporting and with strictly stationary error terms along the lines of the continuous CiC model, and provide a characterization of parallel trends on the latent outcomes within a threshold-crossing model framework: we establish that this type of “Changes-in-Changes” setup holds if and only if a parallel trends-type assumption holds for the functional mean (median) and the scaling parameters. In the binary case and under distributional assumptions, this characterization can nest the commonly used Logit or Probit DiD for binary responses (see Puhani, 2012; Wooldridge, 2023) as these models typically draw on the conventional parallel trends that exclusively affect the mean parameters.
As a second contribution, we derive sharp, nonparametric bounds for the (counter)factual outcome distribution of the treatment group after treatment in the presence of underreporting. We show that, without further restrictions, the bounds on the counterfactual distribution remain uninformative even if an instrumental exclusion restriction for the consumption is available. By contrast, leveraging the CiC structure alongside with an instrumental exclusion restriction and an upper bound on the cumulative misreporting probability, we demonstrate that more informative, tight bounds can potentially be achieved. Our bounds extend the partial identification results of Athey and Imbens (2006) for the discrete outcome case to underreporting. Moreover, the results also complement recent results of Mondal and Wang (2024) for binary outcome models under misreporting by adapting their findings to an ordered choice setup. A suitable estimation and inference procedure for the nonparametric bounds is discussed in the supplementary material.
As a third contribution, we complement the above findings with a set of point identification results for a more structural approach that explicitly models the consumption and reporting behavior, respectively. Specifically, this model generalizes the conventional “false-zero” reporting by allowing survey respondents to report values that are below the actual level, not limited to zero only. We impose restrictions neither on the dependence structure between the unobservables governing the outcome and reporting decisions, nor on their evolution across groups and over time. In this respect, our approach is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically assume the independence between the unobservables (see Hausman et al., 1998; Lewbel, 2000). Our approach also accounts for the potential endogeneity of reporting behavior with respect to treatment status, a frequent concern in observational studies (Arulampalam et al., 2017). We address the identification and estimation challenges posed by the partially observed outcome model by proposing a set of sufficient conditions for the identification of differnt causal parameters of interest for consumption and reporting behavior. Our setup, though akin to the parametric bivariate choice model with fully observed outcomes (e.g., Mourifié and Méango, 2014; Han and Vytlacil, 2017), is notably more complex in its analysis due to the partial observability. We demonstrate that having an exclusion restriction in either the reporting or the consumption equation that exhibits some continuous variation, along with a real analytic condition on the joint distribution of unobservables, is sufficient for point identification of the distributional treatment effects on the treated for consumption or the reporting behavior, respectively. This real analytic assumption, a strong form of smoothness imposed on the copula and the marginal distribution functions of the unobservables, has been used for identification in various contexts including quantile sample selection (Arellano and Bonhomme, 2017), triangular models with endogenous binary treatment (Pereda-Fernández, 2023), and random coefficient discrete choice models (Fox et al., 2012; Fox and Gandhi, 2016).
Finally, as an empirical application, we investigate the impact of recreational marijuana legalization for adults in several U.S. states on the short-term consumption behavior of 8th-grade high-school students. The focus on the last group is particularly relevant as they are widely believed to be most vulnerable to the associated health risks of marijuana consumption. Unlike existing studies which examined the effects of legalization of medicinal or recreational marijuana usage on the mean (cf. Anderson et al., 2015; Wen et al., 2015; Cerdá et al., 2017; Hollingworth et al., 2022), or on the timing of the first (early age) consumption (Williams and Bretteville-Jensen, 2014), we consider the effects of the legalization of recreational marijuana use on the entire consumption distribution of the aforementioned high-school population in the treated states of California, Maine, Nevada, and Massachusetts, thus allowing for the possibility of heterogeneous effects of the reforms. Moreover, in contrast to the existing literature, we address the concerns that consumption behavior may not always be reported truthfully, and that reporting behavior may have been affected by legalization through shifts in stigma, public perception, or general attitudes towards marijuana consumption. Our analysis reveals that, while initial findings without covariates indicate no significant changes, incorporating covariates uncovers a significant decrease in non-consumption probability by around 2 percentage points, and significant increases in consumption probabilities by approximately 1 percentage point at each level. Notably, after accounting for misreporting, the effect sizes increase by more than 50% at each level. Despite these shifts in consumption, our findings do not corroborate significant changes in reporting behavior.
The paper is organized as follows. Section 2 outlines the model setup. Section 3 introduces the concept of distributional parallel trends, providing necessary and sufficient conditions for them to hold within a general threshold-crossing framework. Section 4 extends the framework to account for underreporting, establishing our main identification results. Section 5 empirically investigates the effects of recreational marijuana legalization on the consumption behavior of 8th grade high-school students. Section 6 concludes.
The supplementary material contains auxiliary results for the partial identification model (Section S.1), the proofs of all formal results (Section S.2), an outline of estimation and inference for the nonparametric bounds as well as for the semiparametric approach modeling the reporting behavior explicitly (Section S.3), a Monte Carlo study where we examine the finite sample behavior of the estimators under different sets of exclusion restrictions (Section S.4), additional empirical results (Section S.5). Finally, the supplement also provides a different Copula decomposition as a basis to construct different treatment effects (Section S.6).
2 Setup
In order to introduce the CiC framework for the ordered choice model and the parallel trends characterization in the next section, we start by considering the setup in the absence of underreporting. Let $Y(1)$ and $Y(0)$ denote the potential outcomes for individuals that take on a finite number of $J+1$ values (in a pre-specified period of time) with and without treatment, respectively. In light of the empirical application of Section 5, we will refer to $Y(1)$ and $Y(0)$ as “true” potential consumption levels in the following. For simplicity, we will focus on the case of two groups and two time periods exclusively. The extension to several (pre-treatment) periods is immediate at the cost of more cumbersome notation. We define $G\in\{0,1\}$ as the group membership indicator (where group 1 is the treated group) and $T\in\{0,1\}$ as the time period indicator (where period 1 is after the treatment). The binary treatment indicator for individuals is given by $D=G\cdot T$, so that $D=1$ if and only if $(G,T)=(1,1)$, and $D=0$ otherwise.
Our primary goal lies in identifying distributional treatment effects on the treated (DTTs) for the outcome of interest, such as the effects of marijuana legalization on different levels of drug consumption among adolescents residing in treated states. That is, in the presence of additional covariates $\mathbf{X}$, we aim to identify:
[…]
for different levels $j$ and values $\mathbf{x}$ in the support of $\mathbf{X}$ (conditional on $G=1$ and $T=1$), i.e., values from the covariate distribution in the treated group after the treatment. To distinguish the current focus from alternative treatment effects on reporting which will be introduced in Section 4, we use the superscript “c” to indicate “consumption”. Furthermore, we may also be interested in the unconditional distributional treatment effect on the treated. This can be obtained by “marginalization” with respect to the covariate distribution in the treated group after the treatment:
[…]
where $\mathcal{X}_{11}$ denotes the support of $\mathbf{X}$ given $G=1$ and $T=1$, and ${F}_{\mathbf{X}|D=1}(\cdot)$ the covariate distribution function for the treated group after treatment.
In the absence of underreporting, the first quantity on the right-hand side RHS of (1) is generally nonparametrically identified from the data, while the second term represents a counterfactual distribution that cannot be identified without further assumptions. However, this condition is rarely satisfied when survey data on sensitive behaviors are involved, as such data are often subject to underreporting. Consequently, the presence of underreporting further complicates the identification of both quantities.
Given our primary interest is modeling self-reported ordered outcomes, we adopt a general threshold-crossing model framework for the counterfactual outcomes by introducing a latent outcome variable $Y^{*}(d)$, $d\in\{0,1\}$:
$Y(d)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\kappa_{0}\leq Y^{*}(d)\leq\kappa_{1}\\ 1&\text{ if }\quad\kappa_{1}<Y^{*}(d)\leq\kappa_{2}\\ 2&\text{ if }\quad\kappa_{2}<Y^{*}(d)\leq\kappa_{3},\end{array}\right.$
[…]
3 Parallel Trends for Ordered Outcomes
As is customary in the DiD or the synthetic control literature, we start by modeling the potential outcomes $Y(0)$. To streamline the notation, we adopt the shorthand from Athey and Imbens (2006) in what follows and write:
$Y_{gt}(d)\stackrel{{\scriptstyle}}{{\sim}}Y(d)\mid G=g,T=t$
for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$, where $\sim$ denotes “is distributed as”. We will adopt the same notation for all other random variables as well. Thus, let $\mathbf{X}_{gt}$ denote a vector of covariates for a specific group membership $g$ and time period $t$, and let $\mathcal{X}_{gt}$ and $d_{X_{gt}}$ denote the corresponding support and dimension of that vector, respectively. Then, letting +:=$\mathbb{R}_{+}:=\{x\in\mathbb{R}\mid x>0\}$, we define the latent consumption $Y_{gt}^{\ast}(0)$ to take the following form:
Assumption 1.
For every $(g,t)\in\{0,1\}\times\{0,1\}$:
$Y^{*}_{gt}(0)=\eta_{gt}(\mathbf{X}_{gt})+\lambda_{gt}(\mathbf{X}_{gt})\varepsilon_{gt},$ (2)
where $\eta_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}$ and $\lambda_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}_{+}$ are unknown, real-valued functions, $\varepsilon_{gt}$ is continuously distributed and strictly stationary within each group given $\mathbf{X}_{gt}$.
Here, when one w.l.o.g. normalizes the mean (median) and variance of $\varepsilon_{gt}$ to zero and one, then $\eta_{gt}(\mathbf{X}_{gt})$ can be viewed as the functional mean (median) parameter of the conditional distribution $Y_{gt}^{\ast}(0)\mid\mathbf{X}_{gt}$, and $\lambda_{gt}(\mathbf{X}_{gt})$ as the functional scaling parameter of that distribution. [^4] To clarify the notation, the functions $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$ are assumed to be identical to all units within the group $g$ and in period $t$, while the random variables $Y^{*}_{i,gt}(0)$, $\mathbf{X}_{i,gt}$, $\varepsilon_{i,gt}$ can vary across units. We drop the subscript $i$ to ease the notation. Note that $\eta_{gt}$ and $\lambda_{gt}$ can also be treated as trivial functions if one is interested in unconditional distributions. While the representation of $Y^{*}_{gt}(0)$ in (2) is w.l.o.g. for any given $g\in\{0,1\}$ and $t\in\{0,1\}$ (provided minor regularity conditions on the moments), the strict stationarity on the other hand postulates that the conditional distribution of $\varepsilon_{gt}$ must not change over time within each group. The latter represents a conditional form of Assumption 3.3 in Athey and Imbens (2006) and lies at the heart of the CiC approach, which is particularly valued for its invariance to transformations of the outcome and its intuitive interpretation (e.g., Assunçao et al., 2023; Huber et al., 2022). Finally, note that Assumption 1 encompasses various (parametric and semiparametric) threshold crossing models (Klein and Sherman, 2002; Klein et al., 2015) and is closely related to nonlinear DiD models (Kim and Lee, 2017; Wooldridge, 2023).
We next outline the definition of a CiC model for the latent outcome variable of interest, $Y^{*}_{11}(0)$, which will be used as key identification assumption throughout the paper. For a generic cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$ of a random variable $Y$ with support $\mathcal{Y}$, define its “generalized inverse” by
$F^{-1}:[0,1]\rightarrow\mathbb{R}\cup\left\{-\infty,\infty\right\},\ q\mapsto F^{-1}\left(q\right)$ $:=\inf\left\{y\in\mathcal{Y}\mid q\leq F\left(y\right)\right\},$
where we adopt the convention $\inf\{\emptyset\}=\infty$ and $\inf\{\mathbb{R}\}=-\infty$.
Definition 1.
$Y^{*}_{gt}(0)$ is said to be generated by a CiC model if for every $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, there exists an open interval $I_{\mathbf{x}}\subseteq\mathcal{Y}^{*}_{\mathbf{x},11}(0)\cap\mathcal{Y}^{*}_{\mathbf{x},01}(0)$ such that for all $v\in I_{\mathbf{x}}$:
$F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}^{-1}\left(F_{Y^{*}_{11}(0)\mid\mathbf{X}_{11}}(v\mid\mathbf{x})\mid\mathbf{x}\right)=F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right).$ (3)
We provide two characterizations of this CiC model within our threshold-crossing framework in Proposition 1 below and in Proposition LABEL:PROPS2 of the supplementary material, respectively. The formula in (3) closely resembles the unconditional identification result for the CiC model with continuous outcomes by Athey and Imbens (2006), see Theorem 3.1 of their paper. A similar formula can also be defined for the unconditional counterfactual distribution when the focus shifts to an unconditional analysis.
To illustrate Definition 1 without additional covariates, Figure 1 graphically demonstrates the construction of the counterfactual distribution. Specifically, the counterfactual distribution is derived by adding the difference in the distributions for the control group over time at each level $q^{\prime}$ (corresponding to the quantile $y^{\ast}$) to the treatment distribution function at each level $q$ (corresponding to quantile $y^{\ast}$ in the treatment distribution before treatment).
Note that (3) has often been compared to other “parallel trends”-type assumptions on the CDFs of the counterfactual distributions (see, e.g., Remark 4 in Roth and Sant’Anna, 2023). One such alternative involves imposing a “parallel trends”-type assumption directly on the CDFs of the latent outcome $Y_{gt}^{*}(0)$:
$F_{Y^{*}_{11}(0)}(y)=F_{Y^{*}_{10}(0)}(y)+F_{Y^{*}_{01}(0)}(y)-F_{Y^{*}_{00}(0)}(y)$
for all $y\in\mathbb{R}$. However, this approach introduces undesirable restrictions within some standard location-scale families (see, e.g., Example 2 in Roth and Sant’Anna, 2023), making it less suitable for modeling discrete outcome. For instance, in the family of normal distributions, such a “parallel trends”-type assumption for the CDFs holds only under restrictive conditions - either $Y_{gt}^{*}(0)$ must exhibit strict stationarity over time within each group, or it must remain stable across groups in each period (but not a mixture of these cases or others). This would consequently restrict the (functional) location and scaling parameters $\eta_{gt}$ and $\lambda_{gt}$ to be identical either across groups or over time.
In contrast, our CiC framework for the latent outcome is more flexible and does not impose homogeneity on the model parameters. Moreover, it accommodates standard parametric Probit and Logit DiD models with specified distribution for the error term. [^5] For example, consider a model $Y^{*}(0)=\mathbf{X}^{\prime}\boldsymbol{\eta}+\alpha T+\beta G+\varepsilon$, where $\varepsilon\sim N(0,1)$. It can be readily shown that (3) is satisfied. Given the widespread use of of Probit and Logit DiDs in applied work (e.g. Mayer et al., 2014; Sequeira, 2016; Hoekstra and Sloan, 2022), this flexibility represents a significant advantage of the CiC framework.
Figure 1: Stylized CiC Model
[…]
Note: The figure is a stylized example of a CiC model where the curves in blue denote the marginal distribution functions $F_{Y^{\ast}_{gt}(0)}(\cdot)$ of the counterfactual latent outcome $Y^{\ast}_{gt}(0)$, $g\in\{0,1\}$ and $t\in\{0,1\}$, with the dashed line denoting the imputed function $F_{Y^{\ast}_{11}(0)}(\cdot)$. The curve in red on the other hand denotes the marginal distribution function of the latent outcome $Y_{11}^{\ast}(1)$ under treatment.
Before stating the first identification result, we impose a further (conditional) support condition, which, together with the strict stationarity condition on $\varepsilon_{gt}$, implies $\mathcal{Y}_{\mathbf{x},11}\subseteq\mathcal{Y}_{\mathbf{x},01}$:
Assumption 2.
For every $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$ and and $(g,t)\in\{0,1\}\times\{0,1\}$, $\mathcal{Y}^{*}_{\mathbf{x},11}(0)\subseteq\mathcal{Y}^{*}_{\mathbf{x},10}(0),$ and $\mathrm{int}\left(\mathcal{Y}^{*}_{\mathbf{x},11}(0)\right)$ is nonempty.
The first proposition characterizes the necessary and sufficient condition on the functional parameters for distributional parallel trends to hold:
Proposition 1.
Suppose Assumptions 1 and 2 hold. Then, $Y^{*}_{gt}(0)$ is generated by a CiC model as defined in Definition 1 if and only if
$\eta_{11}(\mathbf{x})-\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}\eta_{10}(\mathbf{x})=\eta_{01}(\mathbf{x})-\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}\eta_{00}(\mathbf{x})\quad\text{and}\quad\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}=\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})},$ (4)
for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$.
Within the general threshold-crossing model given by Assumption 1, Proposition 1 shows that $Y^{*}_{gt}(0)$ is generated by a CiC model if and only if a weighted “parallel trends”-type assumption holds for the functional mean (median) parameters and a proportional “parallel trends”-type holds for the scaling parameters. We note that if time trends exclusively affect the mean (or median) parameters (so $\lambda_{g0}(\mathbf{x})=\lambda_{g1}(\mathbf{x})$), then the results align with the conventional parallel trends, thereby nesting the commonly used Logit or Probit DiD models. More specifically, as can be seen in the first part of (4), the change in the mean (or median) latent outcome for the treated group over time given the covariates, $\eta_{11}(\mathbf{x})-\eta_{10}(\mathbf{x})$, in the absence of treatment, is extrapolated by the change for the control group over the same period, $\eta_{01}(\mathbf{x})-\eta_{00}(\mathbf{x})$, after an appropriate adjustment through the scaling parameters, serving as tailored weighting factors. Here, the weights play the role of correcting the part of the common time trends that cannot be eliminated by the simple mean difference over time within groups. The equality in (4) will serve as a fundamental cornerstone for identifying the counterfactual distributions in Section 4.
Finally, in the supplementary material, we establish another equivalence result to further characterize the class of CiC models in Definition 1. The result helps to clarify the link between the CiC models considered in Athey and Imbens (2006) and our approach.
4 Underreported Outcomes
As emphasized in the introduction, a persistent challenge in analyzing sensitive behaviors, such as the consumption of “social bads”, is the presence of “false zeros” or, more broadly, underreporting. Individuals often conceal their actual participation in such behaviors when responding to retrospective surveys (Nguimkeu et al., 2019). This leads to a partially observable outcome model where a specific outcome value is observed in the data if and only if the actual outcome corresponds to that value and is reported truthfully (cf. Poirier, 1980). To account for this issue, we introduce potential “reported” consumption, denoted by $C(d)$, $d\in\{0,1\}$, which allows for the possibility that $C(d)\neq Y(d)$ with positive probability. The observability rule is then given by $C=D\cdot C(1)+(1-D)\cdot C(0)$.
It is well established that the CiC model with discrete outcomes, even in the absence of misreporting, does not achieve point identification for the counterfactual distribution $F_{Y_{11}(0)}$ (cf. Athey and Imbens, 2006). When underreporting is introduced alongside the discrete nature of the outcome, the identification problem becomes even more complex.
In this section, we address the challenges posed by underreporting in two distinct ways. First, in Subsection 4.1, we derive nonparametric bounds for the counterfactual distribution of the treatment group after treatment by leveraging parallel trends and an exclusion restriction. However, as we will demonstrate, the expressions for the bounds of the counterfactual distribution are only informative when paired with a relatively small upper bound on the cumulative underreporting probability. A different route is taken in Subsection 4.2, where we introduce a semiparametric CiC model for both consumption and reporting decisions, and achieve point identification.
4.1 Partial Identification
In order to derive informative and tight bounds for the potential outcome distribution $F_{Y_{11}(d)}$, $d\in\{0,1\}$, we firstly leverage an instrumental exclusion restriction that does not affect consumption behavior directly.
Assumption 3.
For any $j\in\{0,1\}$ and $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$ and $z\in\mathcal{Z}_{\mathbf{x},gt}$:
$\operatorname{Pr}\left(Y_{gt}(d)\leq j\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)=\operatorname{Pr}\left(Y_{gt}(d)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right).$
An example for an instrumental exclusion $Z_{gt}$ in our context is a “survey cooperation” indicator, which measures respondents’ willingness to cooperate during the survey when consumption behavior is elicited. This indicator is assumed to have no direct effect on consumption $Y_{gt}(0)$, conditional on all other covariates. As is common with instrumental exclusion restrictions, the condition is untestable and the validity of the instrument is subject to the assumptions of the researcher.
Assumption 4.
For any $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$ and $z\in\mathcal{Z}_{\mathbf{x},gt}$:
$10<\operatorname{Pr}\left(C_{gt}(d)\leq 0\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)<\operatorname{Pr}\left(C_{gt}(d)\leq 1\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)<$ 1
Assumption 4 is a non-degeneracy condition on the observed potential outcome $C_{gt}(d)$, $d\in\{0,1\}$, ensuring that all levels $j\in\{0,1,2\}$ lie within the conditional support of reported consumption. Finally, we combine the exclusion restriction from Assumption 3 with a direct bound on the cumulative misreporting probability. The latter becomes necessary because the exclusion restriction alone is insufficient to provide informative bounds for the counterfactual distribution, as demonstrated in Proposition 3 below.
Assumption 5.
For any $j\in\{0,1\}$, $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$:
$\operatorname{Pr}\left(C_{gt}(d)\geq j+1\mid Y_{gt}(d)\leq j,\mathbf{X}_{gt}=\mathbf{x}\right)$ $=0,$
$\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x}\right)$ $\leq\alpha,$
where $\alpha\in[0,1]$ is a known constant.
Assumption 5 explicitly characterizes one-sided misreporting (i.e., underreporting). The first part of Assumption 5 states that all kinds of “overreporting probabilities” are zero. Specifically, if the true potential consumption level is less than or equal to $j$, individuals cannot overstate their consumption to any level greater than $j$ in their reports.
The second part of Assumption 5 introduces an explicit upper bound on the “underreporting probabilities”. This bound can reflect the researcher’s institutional knowledge or serve as a sensitivity parameter. For $\alpha=1$, the assumption imposes no restriction on underreporting, whereas $\alpha=0$ prohibits any underreporting. Note that these assumptions generalize the concept of standard misreporting probabilities for binary outcomes (see, e.g., Mondal and Wang, 2024) to cumulative distribution functions in the context of ordered outcomes.
To state the bounds, we introduce an alternative generalized inverse distribution function, denoted by $F^{(-1)}$ (see Athey and Imbens, 2006). Given a generic cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$ of a random variable $Y$ with support $\mathcal{Y}$, we define
$F^{(-1)}:[0,1]\rightarrow\mathbb{R}\cup\{-\infty,\infty\},\ q\mapsto F^{(-1)}\left(q\right)$ $:=\sup\left\{y\in\mathcal{Y}\mid q\geq F\left(y\right)\right\},$
where we adopt the convention $\sup\{\emptyset\}=-\infty$ and $\sup\{\mathbb{R}\}=\infty$. We obtain the following result:
Proposition 2.
Suppose $Y^{*}_{gt}(0)$ is generated by a CiC model, Assumptions 1-5 hold, and $\mathcal{Y}_{\mathbf{x},gt}=\{0,1,2\}$ for $(g,t)\neq(1,1)$. Suppose $\alpha\geq\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\neq(1,1)}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}$. Then, for each $j\in\{0,1,2\}$ and $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$:
$F_{Y_{11}(1)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[L_{11,\alpha}\left(j\mid\mathbf{x}\right),U_{11}\left(j\mid\mathbf{x}\right)\right],$
$F_{Y_{11}(0)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[L^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right),U^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)\right],$
where
$L^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)$ $:=L_{10,\alpha}\left(\left(U_{00}\right)^{(-1)}\left(L_{01,\alpha}\left(j\mid\mathbf{x}\right)\mid\mathbf{x}\right)\mid\mathbf{x}\right),$
$U^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)$ $:=U_{10}\left(\left(L_{00,\alpha}\right)^{-1}\left(U_{01}\left(j\mid\mathbf{x}\right)\mid\mathbf{x}\right)\mid\mathbf{x}\right),$
and
$L_{gt,\alpha}\left(j\mid\mathbf{x}\right)$ $:=\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-\min\left\{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right),\alpha\right\}}{1-\min\left\{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right),\alpha\right\}},$ (5)
$U_{gt}\left(j\mid\mathbf{x}\right)$ $:=\min_{z\in\mathcal{Z}_{gt,\mathbf{x}}}F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)$ (6)
for $j\in\{0,1\}$, and $L_{gt,\alpha}\left(j\mid\mathbf{x}\right)=U_{gt}\left(j\mid\mathbf{x}\right):=1$ for $j=2$. Moreover, the bounds are sharp when $\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}\leq\alpha<\min_{(g,t)\in\{0,1\}\times\{0,1\}}F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right).$
In addition, if $\alpha<\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}$, then the set of admissible models is empty.
The above proposition derives valid bounds for the conditional cumulative distribution function of the counterfactual $Y_{11}(0)$ and the factual $Y_{11}(1)$. In addition, note that the bounds in Equation (5) and (6) are also valid bounds for the conditional distribution $Y(0)$ of each sub-group with $(g,t)\neq(1,1)$. In fact, when $\alpha=1$ and no bound is placed on underreporting, it is easy to see that
$\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)\in\left[0,\min_{z\in\mathcal{Z}_{gt,\mathbf{x}}}F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)\right]$
for $(g,t)\neq(1,1)$ and $j\in\{0,1\}$, which in turn implies that the bounds for the counterfactual distribution become
$F_{Y_{11}(0)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[0,1\right]$
for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$. On the other hand, observe that Proposition 2 does also not rule out the case of no underreporting when the researcher sets $\alpha=0$. To see this, observe that when $\alpha=0$, it holds that:
$\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)=0$
for all $j\in\{0,1\}$. Under Assumption 3 and the first part of Assumption 4, we can write:
$F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}(j\mid\mathbf{x},z)$
$=$ $\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)$
$\quad+\underbrace{\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)}_{=0}\times\left(1-\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)\right).$
However, under Assumption 3 and $\alpha=0$, this equality holds if and only if $F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}(j\mid\mathbf{x},z)=F_{C_{gt}\mid\mathbf{X}_{gt}}(j\mid\mathbf{x})$ for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$ and $j\in\{0,1\}$, which in turn restricts the set of admissible, observable distributions. [^6] Note that in this case $\alpha=\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}=0$ for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$. Finally, in Section S.3.1 of the supplementary material, we comment on the possibility to estimate these bounds nonparametrically and to carry out inference via the nonparametric i.i.d. bootstrap following the ideas of Fang and Santos (2019) and Poirier and Masten (2020).
4.2 Point Identification
As the expressions for the bounds of the counterfactual distribution $Y_{11}(0)$ are only informative when paired with a relatively small upper bound on the cumulative underreporting probability (even the bounds may still be wide in practice), we take a different route in this subsection. Specifically, we introduce a semiparametric CiC model that jointly accounts for both consumption and reporting decisions. While this approach is more restrictive, as it imposes additional structure on the model, it will allow us to define and to identify new treatment parameters of interest, such as the DTT on the reporting intention.
To formalize the approach, let $R(d)$, $d\in\{0,1\}$ denote the potential “reporting intention”, analogous to $Y(d)$, with $R(d)$ taking on $J+1$ discrete value values. Here, we interpret $R(d)$ as the (potential) level up to which an individual commits to reporting her consumption truthfully. The potential outcome of interest, $C(d)$, $d=0,1$, is then given by
$C(d)=\min\left\{Y(d),R(d)\right\}.$ (7)
Thus, respondents will truthfully report their consumption $Y(d)$ if and only if it is less than or equal to the reporting intention $R(d)$. For instance, if $Y(0)=2$ and $R(0)=1$, then, in the absence of treatment, the respondent would only report consumption at level 1 even though the actual level is 2. It is worth noting that this formulation of underreporting behavior in (7) allows for more flexibility than standard “false zero” reporting setups, where respondents may underreport a positive outcome level exclusively as a “false zero” (see Brown et al., 2018; Greene et al., 2018; Nguimkeu et al., 2019). The above definition also bears similarities with the binary selection setup in Klein et al. (2015). However, additional complications in the identification arise due to the partial observability of the outcome of interest, stemming from the unknown reporting decision (see also Poirier (1980)). That is, unlike a selection setup where the selection indicator is fully observed, $R(d)$ can be only partially inferred for each $d$ from the observed joint outcome of individual consumption and reporting decisions. Importantly, also note that in our setup reporting behavior is allowed to change in response to the treatment status $d$. For example, changes in the legalization of recreational marijuana usage for adults may alter the perceived harmfulness or stigma associated with its consumption, thereby also affecting individuals’ willingness to report truthfully.
The introduction of $R(d)$ allows us to define new policy parameter of interest such as the for the treated on the reporting intention:
$\tau^{r}(j\mid\mathbf{z}):=$ $\Pr\left(R(1)=j\mid D=1,\mathbf{Z}=\mathbf{z}\right)-\Pr\left(R(0)=j\mid D=1,\mathbf{Z}=\mathbf{z}\right),$ (8)
for different levels $j$ and values $\mathbf{z}$ in the support of $\mathbf{Z}$ given $G=1$ and $T=1$, where $\mathbf{Z}$ denotes a covariate vector nonnested within $\mathbf{X}$. The parameter in (8) provides insights into the possible changes in the probabilities of reporting on individuals in response to the treatment. This may also shed light on how factors like stigma and shame influence the reporting behavior. In analogy to $Y^{\ast}(0)$, we conclude by defining the latent reporting intention variable $R^{\ast}(0)$:
$R(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\iota_{0}\leq R^{*}(0)\leq\iota_{1}\\ 1&\text{ if }\quad\iota_{1}<R^{*}(0)\leq\iota_{2}\\ 2&\text{ if }\quad\iota_{2}<R^{*}(0)\leq\iota_{3},\end{array}\right.$ (12)
where $\iota_{0}<\iota_{1}<\iota_{2}<\iota_{3}$, setting again $\iota_{0}=-\infty$ and $\iota_{3}=\infty$, and (w.l.o.g.) $\iota_{1}=0$ and $\iota_{2}=1$.
4.2.1 Counterfactual Distribution
Building upon the observation rule outlined in (7) and the latent outcome formulation of $R(0)$ in (12), we consider semiparametric versions of our model for $Y_{gt}^{\ast}(0)$, where the distribution of $\varepsilon_{gt}$ remains unspecified and further structure is placed on the functional form of $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$, respectively. Specifically, for every $(g,t)\in\{0,1\}$, we assume that $\eta_{gt}(\mathbf{X}_{gt})=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ with $\overline{\mathbf{X}}_{gt}=(1,\mathbf{X}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\eta}}_{gt}=(\eta_{c,gt},\boldsymbol{\eta}_{gt}^{\prime})^{\prime}$, while $\lambda_{gt}(\mathbf{X}_{gt})=\lambda_{gt}>0$. Thus:
$Y_{gt}^{\ast}(0)=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}+\lambda_{gt}\varepsilon_{gt}.$ (13)
Equation (13) is a standard functional specification for ordered choice models (Klein and Sherman, 2002), and allows the mean of $Y_{gt}^{\ast}(0)$ to vary with the covariates $\mathbf{X}_{gt}$, but not the scaling parameter. Similar to the structure of $Y_{gt}^{\ast}(0)$, we also model $R_{gt}^{\ast}(0)$ as a function of additional covariates $\mathbf{Z}_{gt}$:
$R_{gt}^{\ast}(0)=\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}+\zeta_{gt}\nu_{gt},$ (14)
where $\overline{\mathbf{Z}}_{gt}=(1,\mathbf{Z}_{gt}^{\prime})^{\prime}$, $\overline{\boldsymbol{\pi}}_{gt}=(\pi_{c,gt},\boldsymbol{\pi}_{gt}^{\prime})^{\prime}$ and $\zeta_{gt}>0$. The covariate vectors $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ are allowed to share common elements. Depending on the treatment parameter of interest, however, it will be required that one vector includes a distinct continuous variable not present in the other.
Under Assumption 6 below, for each $j\in\{0,1,2\}$, the upper tail probability of consumption in the absence of treatment can be expressed as:
$\Pr\left(C_{gt}(0)\geq j\mid\mathbf{X}_{gt},\mathbf{Z}_{gt}\right)$ $=\Pr\left(Y_{gt}(0)\geq j,R_{gt}(0)\geq j\mid\mathbf{X}_{gt},\mathbf{Z}_{gt}\right)$
$=\Pr\left(\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}+\lambda_{gt}\varepsilon_{gt}\geq\kappa_{j},\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}+\zeta_{gt}\nu_{gt}\geq\iota_{j}\right)$ (15)
$=\overline{C}_{gt}\left(F_{-{\varepsilon_{gt}}}\left(\frac{\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}}{\lambda_{gt}}\right),F_{-{\nu}_{gt}}\left(\frac{\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}}{\zeta_{gt}}\right)\right).$
where $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$ denote CDFs of $-\varepsilon_{gt}$ and $-\nu_{gt}$, respectively, and $\overline{C}_{gt}(\cdot,\cdot)$ denotes a nonparametric Copula function. As noted in Section 2, the above setup bears similarities with the distributional sample selection model proposed by Chernozhukov et al. (2023), as well as the bivariate Probit model with a binary endogenous regressor considered by Han and Vytlacil (2017). However, a crucial distinction of our model, represented by (15), lies in the fact that the reporting intention $R_{gt}(0)$ is not directly observable. As a result, the second argument in the stated expression cannot be separately identified and estimated. This suggests that the identification results presented in the aforementioned papers are not directly applicable to our context. This insight is supported by the simulation results in Section LABEL:sec:simulation of the supplementary material. These simulations suggest that parameters from the model can, in some cases, still be poorly estimated even if one exclusion restriction is in place.
In what follows, let $X_{1,gt}$ ($Z_{1,gt}$) denote the first element of the vector $\mathbf{X}_{gt}$ ($\mathbf{Z}_{gt}$), and $\mathbf{X}_{-1,gt}$ ($\mathbf{Z}_{-1,gt}$) the vector containing the remaining ones. We make the following assumption:
Assumption 6.
(i) $-\varepsilon_{gt}$ and $-\nu_{gt}$ are strictly stationary within each group with zero median and unit variance, jointly independent of $(\mathbf{X}_{gt}^{\prime},\mathbf{Z}_{gt}^{\prime})^{\prime}$, and have unknown, real analytic and strictly monotonic CDFs $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$, respectively.
(ii) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, the copula function $\overline{C}_{gt}(\cdot,\cdot)$ defined in (15) is strictly monotonic and real analytic in each of its coordinates.
Assumption 7.
(i) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, (a) the variable $X_{1,gt}$ is excluded from the vector $\mathbf{Z}_{gt}$; (b) the variable $Z_{1,gt}$ is excluded from the vector $\mathbf{X}_{gt}$.
(ii) For each $(g,t)\neq(1,1)$, $\eta_{1,gt}\neq 0$ and $\pi_{1,gt}\neq 0$ and there exists some Cartesian product of two open intervals $I_{1}\times I_{2}\subseteq\mathbb{R}\times\mathbb{R}$ such that
$\operatorname{Pr}\left(\left(X_{1,gt},Z_{1,gt}\right)\in I_{1}\times I_{2}\mid\overline{\mathbf{W}}_{gt}\right)>0$
with probability one, where $\overline{\mathbf{W}}_{gt}=\left(1,\mathbf{X}_{-1gt}^{\prime},\mathbf{W}_{-1gt}^{\prime}\right)^{\prime}$.
(iii) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\mathcal{\overline{X}}_{gt}$ and $\mathcal{\overline{Z}}_{gt}$ are not contained in a proper linear subspace of $\mathbb{R}^{d_{X_{gt}}+1}$ and $\mathbb{R}^{d_{Z_{gt}}+1}$, respectively, where $d_{X_{gt}}$ and $d_{Z_{gt}}$ denote the dimension of the vectors $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$.
While the covariate independence in Assumption 6(i) is restrictive, it is important to note that we do not impose any constraints on the correlation between the unobservables $\varepsilon_{gt}$ and $\nu_{gt}$. This allowance caters to scenarios where certain types of “consumers” may be more or less likely to misreport than others. This aspect of our model is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically imposes (conditional) independence between the unobservables governing the outcome and reporting decisions (see Hausman et al., 1998; Lewbel, 2000). On the other hand, the other parts of Assumption 6 impose further conditions on the joint and marginal distributions of the unobservables. In particular, the real analytic property will allow for identification of the counterfactual distribution based on an extrapolation argument, which has also been featured in the microeconometric (see Arellano and Bonhomme, 2017; Pereda-Fernández, 2023) and the Industrial Organization identification literature (see Iaria and Wang, 2022, and references therein). Specifically, the extrapolation is based on the fact that if two real analytic functions coincide on an open neighborhood, then they coincide everywhere, and in particular also in the limit (if it exists). Identification based on such extrapolation is in contrast to “identification at infinity” type arguments, which require the existence of excluded (instrumental) variables with a large support (e.g., Chamberlain, 1986; Lewbel, 2007; Corradi and Gutknecht, 2023). Moreover, note that the real analytic property is satisfied by several well-known Copula families such as the Gaussian, Frank, Clayton, Plackett, Joe, or Gumbel Copula families. Indeed, while Assumptions 7 (i)-(a) as well as (i)-(b) postulate an exclusion restriction for the identification of the parameters of the consumption and the reporting decision, respectively, part (ii) is rather “mild” in its continuity requirement. [^7] Note that Assumption 7(i)-(b) together with Assumption 6(i) imply Assumption 3 from Subsection 4.1. Specifically, it requires one continuous element in each vector, one of which is the instrumental variable. Finally, note that under the assumption that both $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ contain an excluded, continuous variable, we may actually not only identify counterfactual parameters for either the consumption or the reporting decision as in Proposition 3 below, but for both simultaneously. We obtain the following identification result:
Proposition 3.
Suppose Assumptions 6 holds and consider the models in (13)-(14). Also, suppose that the CiC model from Definition 1 holds for $Y_{gt}^{*}(0)$ as well as for $R_{gt}^{*}(0)$, and that $\mathcal{X}_{11}\subseteq_{(g,t)\neq(1,1)}\mathcal{X}_{gt}$.
(i) Under Assumption 7 (i)-(a), for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\lambda_{gt}$, and $F_{-\varepsilon_{gt}}(v)$ are identified. In particular, we have that:
$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}},$ (16)
(ii) Under Assumption 7 (i)-(b), for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\pi}}_{gt}$, $\zeta_{gt}$, and $F_{-\nu_{gt}}(v)$ are identified. In particular, we have that:
$\overline{\boldsymbol{\pi}}_{11}=\overline{\boldsymbol{\pi}}_{01}+\frac{\overline{\boldsymbol{\pi}}_{10}-\overline{\boldsymbol{\pi}}_{00}}{\zeta_{00}/\zeta_{01}}\quad\text{and}\quad\zeta_{11}=\frac{\zeta_{01}\zeta_{10}}{\zeta_{00}}.$ (17)
It should be emphasized that the CiC conditions solely set constraints on the change of the marginal distributions of counterfactual outcomes over time. However, these assumptions do not place any constraints on the evolution of the dependence structure between the outcome and reporting decisions. Additionally, it is noteworthy that the identification result presented in Proposition 3 requires knowledge of neither the marginal distributions nor the Copula function, thereby classifying it as semiparametric. Nevertheless, when transitioning to estimation, one has to select specific parametric forms for the Copula function $\overline{C}_{gt}(\cdot,\cdot)$ as well as the distributions $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$. This is because the identification based on the real analytic assumption is not “constructive” and does not provide a direct pathway to construct estimators.
4.2.2 Distributional Treatment Effects
To identify the different DTTs, we start by setting up a model for $Y_{11}(1)$ and $R_{11}(1)$, respectively. As modeling $Y_{11}(1)$ become indispensable with misreporting, we postulate an analogous model structure for $Y_{11}(1)$ as for $Y_{11}(0)$ without misreporting. Specifically, drawing on the model for $Y_{11}^{\ast}(0)$ in (13), we define the latent consumption in the presence of treatment as:
$Y_{11}^{\ast}(1)=\overline{\mathbf{X}}_{11}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}+\lambda_{11}^{(1)}\varepsilon_{11}^{(1)}.$ (18)
where we introduced the superscript (1) to distinguish the parameters of the potential outcome under treatment from those under control. From this point forward, we will accordingly add the superscript $(0)$ to parameters in the absence of treatment. Drawing on the model for $R_{11}^{\ast}(0)$ in (14), we define analogously:
$R_{11}^{\ast}(1)=\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}+\zeta_{11}^{(1)}\nu_{11}^{(1)},$ (19)
and adopt the following assumption:
Assumption 8.
$-\varepsilon_{11}^{(1)}$ and $-\nu_{11}^{(1)}$ are strictly stationary within each group with zero median and unit variance, jointly independent of $(\mathbf{X}_{11}^{\prime},\mathbf{Z}_{11}^{\prime})^{\prime}$, and have unknown, real analytic and strictly monotonic CDFs $F_{-\varepsilon_{11}^{(1)}}$ and $F_{-\nu_{11}^{(1)}}$, respectively. Also, the copula function $\overline{C}_{11}^{(1)}(\cdot,\cdot)$ defined below is strictly monotonic and real analytic in each of its coordinates:
$\Pr\left(C_{11}(1)\geq j\mid\mathbf{X}_{11},\mathbf{Z}_{11}\right)$ $=\overline{C}_{11}^{(1)}\left(F_{-{\varepsilon_{11}^{(1)}}}\left(\frac{\overline{\mathbf{X}}_{11}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j}}{\lambda_{11}^{(1)}}\right),F_{-{\nu}_{11}^{(1)}}\left(\frac{\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}-\iota_{j}}{\zeta_{11}^{(1)}}\right)\right).$
Finally, $\eta_{1,11}^{(1)}\neq 0$ and $\pi_{1,11}^{(1)}\neq 0$.
We have the following result for identification of the conditional DTTs for consumption and reporting, respectively, in the presence of misreporting:
Proposition 4.
Suppose Assumptions 6, 7, and 8 hold, and consider the model in (13)-(14) and (18)-(19). Also, suppose the CiC model from Definition 1 holds for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$. Then, for all $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$:
(i) Under Assumption 7 (i)-(a), $\tau^{c}(j)$ is identified as:
$\tau^{c}(j\mid{\mathbf{x}})=$ $\left(F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j+1}}{\lambda^{(1)}_{11}}\right)-F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j}}{\lambda^{(1)}_{11}}\right)\right)$
$-\left(F_{-\varepsilon_{11}^{(0)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}-\kappa_{j+1}}{\lambda_{11}^{(0)}}\right)-F_{-\varepsilon_{11}^{(0)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}-\kappa_{j+1}}{\lambda_{11}^{(0)}}\right)\right).$
(ii) Under Assumption 7 (i)-(b), $\tau^{r}(j)$ is identified as:
$\tau^{r}(j\mid{\mathbf{z}})=$ $\left(F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j+1}}{\zeta^{(1)}_{11}}\right)-F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j}}{\zeta^{(1)}_{11}}\right)\right)$
$-\left(F_{-\nu_{11}^{(0)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(0)}_{11}-\iota_{j+1}}{\zeta_{11}^{(0)}}\right)-F_{-\nu_{11}^{(0)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(0)}_{11}-\iota_{j}}{\zeta_{11}^{(0)}}\right)\right).$
Interest may also extend to unconditional DTTs, which are straightforward to identify and to construct. Additionally, introducing stability conditions to the copula function can help identify the counterfactual copula function. This aspect is relevant for understanding the impact of treatment on the dependence structure between individual consumption decisions and reporting intentions. In Section LABEL:sec:copuladecomp of the supplementary material, we present a decomposition of another possible causal quantity. This effect is characterized by the difference between potential self-reported outcome distributions under treatment and control, where the counterfactual copula function is involved.
Remark 1: As the identification argument presented in this subsection is not constructive, estimation may be carried out via Maximum Likelihood (ML) relying on a specific choice for the copula function. In Section S.3.2 of the supplementary material, we present corresponding ML estimators for panel as well as repeated cross-sectional data, and discuss their asymptotic properties. Specifically, as all model parameters are estimated jointly, note that consistent estimation requires both Assumption 7 (i)-(a) and (i)-(b) to hold simultaneously. For a given choice of the marginal distribution $F_{-\varepsilon}$ and $F_{-\nu}$, respectively, the ML estimators of the conditional DTTs outlined in Proposition 4 are given by:
$\widehat{\tau}^{c}(j\mid\mathbf{x})$ $:=\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$
$\widehat{\tau}^{r}(j\mid\mathbf{z})$ $:=\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(1)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j}}{\widehat{\zeta}^{(1)}_{11}}\right)\right]-\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(0)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j}}{\widehat{\zeta}^{(0)}_{11}}\right)\right]$
for $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$, where
$\widehat{\overline{\boldsymbol{\eta}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\eta}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\eta}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\eta}}}_{00}^{(0)}}{\widehat{\lambda}_{00}^{(0)}/\widehat{\lambda}_{01}^{(0)}},\quad\widehat{\lambda}_{11}^{(0)}:=\frac{\widehat{\lambda}_{01}^{(0)}\widehat{\lambda}_{10}^{(0)}}{\widehat{\lambda}_{00}^{(0)}},\quad\widehat{\overline{\boldsymbol{\pi}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\pi}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\pi}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\pi}}}_{00}^{(0)}}{\widehat{\zeta}_{00}^{(0)}/\widehat{\zeta}_{01}^{(0)}}\quad\text{and}\quad\widehat{\zeta}_{11}^{(0)}:=\frac{\widehat{\zeta}_{01}^{(0)}\widehat{\zeta}_{10}^{(0)}}{\widehat{\zeta}_{00}^{(0)}}$
are the corresponding ML estimates. Similarly, the construction of “marginalized” DTTs for the treated group after treatment (see, e.g., Chernozhukov et al., 2013) are:
[…]
Here, $\widehat{F}_{\mathbf{X}_{11}}(\cdot)$ represents an estimator of the covariate distribution.
Remark 2: While Proposition 2 does not allow any conclusions about the relationship of $\overline{\boldsymbol{\eta}}_{gt}$ and $\lambda_{gt}$ across pre-treatment periods, note that the ML estimator mentioned in previous remark can be used to conduct “pre-treatment” tests of the CiC model structure in analogy to the popular “pre-trend” test. That is, using the consumption equation as illustration, we may test the implied CiC relationship in Equation (11) from Proposition 2 in pre-treatment periods $t=0,-1,-2,\ldots$. Specifically, under the assumption that Equation (11) holds also for any given pre-treatment period $t\in\{0,-1,-2,\ldots\}$, we may test:
$H_{0}:\overline{\boldsymbol{\eta}}^{(0)}_{1t}-\overline{\boldsymbol{\eta}}^{(0)}_{0t}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}\left(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)}\right)=\mathbf{0}\quad\text{and}\quad\frac{\lambda^{(0)}_{1t}}{\lambda^{(0)}_{1(t-1)}}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}=0,$
against
$H_{A}:\overline{\boldsymbol{\eta}}^{(0)}_{1t}-\overline{\boldsymbol{\eta}}^{(0)}_{0t}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}\left(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)}\right)=\mathbf{0}\quad\text{and/ or}\quad\frac{\lambda^{(0)}_{1t}}{\lambda^{(0)}_{1(t-1)}}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}=0.$
via a Likelihood Ratio test comparing an unrestricted log-likelihood as in Equation (15) with a restricted version that satisfies $\overline{\boldsymbol{\eta}}^{(0)}_{1t}=\overline{\boldsymbol{\eta}}^{(0)}_{0t}+\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)})$ and $\lambda^{(0)}_{1t}=\frac{\lambda^{(0)}_{0t}\lambda^{(0)}_{1(t-1)}}{\lambda^{(0)}_{0(t-1)}}$, respectively.
5 Empirical Application
The intense debate surrounding the legalization of marijuana (cannabis) and its impact on individual consumption behavior, and more generally, on society, has been a pivotal policy issue in Western countries for decades. In many of these nations, the possession and consumption of marijuana are illegal, primarily due to concerns over its contribution to crime escalation and the associated societal and economic costs (Brinkman and Mok-Lamme, 2019). Additionally, there are health-related concerns, with evidence suggesting that early-age cannabis consumption can lead to lower educational attainment (Van Ours and Williams, 2009; Moore et al., 2007; Macleod et al., 2004) and cognitive deficits in verbal learning and memory tasks (Solowij et al., 2011). The potential of cannabis serving as a gateway to more harmful drug use has also been documented (see Pudney, 2010, and references therein).
In this section, we investigate the effects of recreational marijuana legalization for adults in several U.S. states on the consumption behavior of 8th-grade high-school students in those states in 2016. The focus on these minors, typically aged 13 to 14, is critical due to their heightened vulnerability to the health risks associated with marijuana consumption. Unlike existing studies, which predominantly examined legalization effects on average usage or consumption (see Anderson et al., 2015; Wen et al., 2015; Cerdá et al., 2017; Hollingworth et al., 2022), or on the timing of the first (early age) consumption (Williams and Bretteville-Jensen, 2014), we investigate the distributional effects of recreational marijuana legalization on the entire consumption distribution within the aforementioned high-school cohort in various states. [^8] Unlike Hollingworth et al. (2022), who consider mean effects for the legalization of recreational marijuana usage on underage adolescents among others, the survey data used in this analysis exclusively covers enrolled high-school students (see next paragraph). Thus, it does not provide insights into effects on high-school drop-outs. Additionally, we account for the possibility that consumption behavior may not always be reported truthfully, and more importantly, that reporting behavior itself may be affected by the legalization, possibly due to shifts in stigma, public perception, or attitudes towards marijuana consumption.
The data is composed of several cross-sectional waves of the “Monitoring the Future” survey, an annual survey conducted in the United States gathering data on attitudes, behaviors, and values of American adolescents currently enrolled in high school. [^9] The “Monitoring the Future” survey is funded by the National Institute on Drug Abuse (NIDA). For more information, see Monitoring the Future, Institute for Social Research, University of Michigan (https://monitoringthefuture.org/). Specifically, the survey for 8th graders, which is conducted annually, covers a wide range of topics including substance (ab)use such as the consumption of marijuana and variants thereof. The primary sample of our empirical analysis consists of 46,472 high-school students, sampled as repeated cross-sections from 2015 to 2018 across all contiguous U.S. states except Oregon, Washington, and Colorado. With this sample, our focus is on the recreational marijuana legalization for adults aged 21 or older in California, Nevada, Massachusetts, and Maine in November 2016, examining its impact on 8th-grade students’ (short-term) consumption behavior. [^10] These states had previously legalized medicinal marijuana use before the observation period. Moreover, even though private cultivation of marijuana for personal consumption became legal (almost) immediately, the first non-medical retail sales of marijuana did not start until some time later in most cases, (e.g. not before 2018 in Maine and California). We classify the observations from 2015-2016 into the pre-treatment period ($t=0$), and those from 2017-2018 into post-treatment period ($t=1$). On the other hand, we group the observations from California, Nevada, Massachusetts and Maine into the treated group ($g=1$), and the rest into the control group ($g=0$). This state grouping is mainly owed to data privacy rules imposed by the hosting institute, preventing a more disaggregate analysis. It is also important to note that no other major marijuana-related legalization took place in the U.S. during this observation period, either for medicinal or recreational purposes. Finally, given these surveys were conducted in springs, we consider the 2016 data to be sufficiently removed from the actual legalization to avoid major anticipation effects.
The outcome variable of interest (cons_30days), based on the survey responses to the question “On how many occasions have you used marijuana (weed, pot) or hashish (hash, hash oil) … in the last 30 days?”, is categorized into three levels: “0” (never used), “1” (used 1-2 times), and “2” (used more than 2 times).  [^11] Table LABEL:Summary1 in Section LABEL:sec:additionalempirical of the supplement presents descriptive statistics of the outcome variable and other controls. Similar to Hollingworth et al. (2022), we not only include individual-level covariates like gender, age, ethnicity, and parental education, but also certain state-level control variables such as unemployment rate and median income, fixed at pre-treatment levels. [^12] State-level unemployment rates as well as the state median income were constructed from data of the Bureau of Labor Statistics (https://www.bls.gov/). Finally, we employ state-level marijuana prices from 2015 as an instrument for the consumption (state_price_m), and a “survey cooperation” measure (trust) for the reporting, following Greene et al. (2018) and Brown et al. (2018). [^13] Marijuana price data source:
https://oxfordtreatment.com/substance-abuse/marijuana/average-cost-of-marijuana/ (last accessed: 12-10-2023).
Here, marijuana prices are presumed to exclusively influence consumption decisions. The plausibility of the exclusion restriction relies on the assumption that consumption behavior is sensitive to prices, while reporting behavior is not, given that we also control for other state-level variables. On the other hand, “survey cooperation”, defined as the proportion of missing survey responses (based solely on questions common to all questionnaire forms and not directly related to drug use), is assumed to provide an indication of the respondents’ willingness to cooperate in the survey, unrelated to marijuana consumption itself, as the reporting decision occurs post-consumption only.
In Table 1, we present our initial findings for a model without underreporting and without additional control variables. Our first set of results (gdid) suggests that $\tau^{c}_{11}(j)$ is not statistically different from zero for $j\in\{0,1,2\}$ at any conventional levels. These results align with the estimates from the more restrictive ordered (pdid_ordered) and binary choice (pdid_probit) DiD models, which include the binary time period and group indicators and their interaction as regressors.
Table 1: Unconditional Treatment Effects (No Misreporting)
pdid pdid (sd) gdid_ordered gdid_ordered (sd) gdid gdid (sd)
$\tau^{c}_{11}(0)$0.0071 0.0056 -0.0067 0.0056 -0.0071 0.0056
$\tau^{c}_{11}(1)$ 0.0071 0.0056 0.0027 0.0045 0.0047 0.0040
$\tau^{c}_{11}(2)$ NA NA 0.0039 0.0033 0.0025 0.0042
Notes: Data between 2015-2018 are used.  The table contains the estimated distributional treatment effects $\tau_{11}^{c}(j)$, $j\in\{0,1,2\}$ for the Probit DiD with binary outcomes (pdid_binary), the Probit DiD with ordered outcomes (pdid_ordered), and generalized DiD proposed in this paper (gdid). For the Probit DiD with binary outcomes, $\tau^{c}_{11}(1)$ indicates the effect for users with outcomes $\geq$ 1. Standard errors are computed via the Delta method and can be found in the corresponding columns next to the coefficient estimates.
Next, we move to specifications without misreporting but including covariates to account for differences across groups in terms of observed characteristics. We present estimates of $\tau_{11}^{c}(j)$ in Table 2 (first panel), directing readers to Section LABEL:sec:additionalempirical in the supplementary material for parameter estimates. Given nearly identical variance parameters across groups and time in Table 1, we constrain $\lambda_{gt}$ to be constant in this specification. [^14] Unreported results (available upon request) show that allowing for flexible variance parameters does not alter the estimates qualitatively. Notably, the estimates of $\widehat{\tau}_{11}^{c}(j)$ now appear to be larger in absolute terms, with a significant decrease in non-consumption probability by around 2 percentage points, and significant increases in consumption probabilities by approximately 1 percentage point. Compared to the unconditional results, conditioning on additional regressors seems to amplify the effect sizes in absolute terms.
In the final step, we estimate models with covariates incorporating misreporting according to Section 4 (see second panel of Table 2). As before, we only present estimates of the marginalized treatment effects $\tau_{11}^{c}(j)$ and $\tau_{11}^{r}(j)$ on consumption and misreporting, respectively, as defined in (20), with further details available in Section LABEL:sec:additionalempirical of the supplement. It is noteworthy to mention that the slope estimates for the consumption decision closely mirror those from the specification without misreporting. Both of the instruments are found to be significant at the 5 or 10% level across groups and time (except for the treatment group after treatment). [^15] The positive and significant effects for marijuana prices align with other studies (e.g. Greene et al., 2018), potentially reflecting drug quality and availability effects, in the absence of quality as a control variable). Notably, we see more pronounced effect sizes after accounting for misreporting (increased by more than 50% for all levels), with all $\widehat{\tau}_{11}^{c}(j)$ being significant at the 5% level. No significant treatment effects are found on the reporting behavior. These findings are robust to a change in the Copula function to the asymmetric Clayton Copula.
Table 2: Distributional Treatment Effects (Marginalized)
Without Misreporting
gdid gdid (sd)
$\tau_{11}^{c}(0)$0.0200 0.0082
$\tau_{11}^{c}(1)$ 0.0088 0.0036
$\tau_{11}^{c}(2)$ 0.0113 0.0047
With Misreporting
gdid_frank gdid_frank (sd) gdid_clayton gdid_clayton (sd)
$\tau^{c}_{11}(0)$0.0327 0.0137 -0.0327 0.0141
$\tau^{c}_{11}(1)$ 0.0138 0.0057 0.0138 0.0058
$\tau^{c}_{11}(2)$ 0.0189 0.0081 0.0189 0.0083
$\tau^{r}_{11}(0)$ 0.1199 0.1213 0.1242 0.1442
$\tau^{r}_{11}(1)$ 0.0125 0.0116 0.0179 0.0268
$\tau^{r}_{11}(2)$0.1324 0.1307 -0.1421 0.1698
Data between 2015-2018 are used.  The first panel of results (“Without Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$, $j\in\{0,1,2\}$, for the model estimated without misreporting. The treatment effect estimates have been marginalized with respect to the distribution of $\mathbf{X}_{11}$. The second panel of results (“With Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$ and $\tau_{11}^{r}(j)$, $j\in\{0,1,2\}$, for the model with misreporting. The treatment effect estimates have also been marginalized with respect to the distribution of $\mathbf{X}_{11}$. The first two columns display results for a Frank Copula specification, the last two columns show corresponding results for a Clayton Copula specification. Standard errors are computed via the Delta method and can be found in the corresponding columns next to the coefficient estimates.
In summary, the empirical findings of this section reveal nuanced effects of recreational marijuana legalization on 8th-grade high-school students’ short-term consumption and reporting behavior. Our initial analysis without misreporting and covariates showed no significant changes in consumption probabilities. However, after introducing covariates, we observed a more pronounced impact, with a significant decrease in non-consumption probability and a significant increase in consumption probabilities. These effect sizes were further amplified and remain significant after accounting for misreporting, while no significant treatment effects are found on the reporting behavior. Nevertheless, it should be noted that, as a limitation of our study, the categorization of the outcome variable focuses on usage frequency and does not capture the doses consumed per occasion, potentially overlooking variations in consumption intensity.
6 Conclusion
This paper develops a Difference-in-Differences model for discrete, possibly underreported ordered outcomes in repeated cross-sectional or panel data. The models are based on the continuous CiC model of Athey and Imbens (2006). We extend this CiC model for discrete outcomes to partially observable outcomes. Special features of this extended model are that it allows the underreporting behavior to be correlated with the outcome of interest and that underreporting may depend on the treatment status itself. We provide point identification conditions for a semiparametric version of the model based on a real analytic extrapolation assumption, and suggest a flexible estimator for a parametric version of the model.
We apply our methodology to investigate the distributional effects of the legalization of recreational marijuana usage for adults in the U.S. states of California, Nevada, Massachusetts, and Maine in November 2016 on the consumption behavior of 8th-grade high-school students. Our findings suggest that there are small, positive significant effects on the (short-term) consumption probabilities, but no significant changes in the reporting behavior as a result of the legalization. Importantly, correcting for misreporting increases effect sizes by more than 50%.